Ticket Number: HEA-NOC/20030218-01 Ticket Status: UPDATE
Ticket Type: Scheduled Ticket Source: HEA-NOC
Ticket Scope: Site Site/Line: Teleute
Ticket Owner: [log in to unmask] Problem Fixer: HEA-NOC
Ticket Opened: Problem Start: 20030219 09:30 UTC
Ticket Closed: Problem End: 20030219 10:00 UTC
OSPF packet size errors from Teleute.
RCSI will experience a short outage of approximately 5 minutes. The other circuits on Teleute are all backups so will not be service affecting.
2003 0218 11:00
ATM card has been RMA'd and will be replaced on Wednesday 18th Feb. at 9:30 A.M. by NOC duty engineer DC. All backup ATM circuits will experience a short downtime.
This ticket should read Wednesday 19th Feb.
20030219 09:55 UTC - DC - Replaced card. All ATM circuits
passing traffic normally. Leaving ticket open for 24 hours
20030226 10:50 UTC - JL
This packet size error is still occuring, Lancomms requested a sh tech from
2 customer routers whose ATM subints were causing the errors. This has been done so currently awaiting an update from Lancomms.
20030227 13:23 UTC - JL
Lancomms have informed us that the TAC engineer has passed this problem to cisco development, awaiting response.
20030307 16:15 UTC - JL
Cisco requested a detailed dump of IP packets on Teleute. This has been sent to the TAC and we are awaiting their response.
20030417 14:50 UTC +1 - JL
Lancomms and cisco TAC have advised rolling back to 12.2(2)T2. IOS rollback will take place during maintenance window on Wed 23rd April at 9:00 a.m. UTC +1
20030428 11:40 UTC +1 - JL
The new IOS image 12.2(2)T2 has been running on Teleute since Wednesday April 23rd and no errors have been reported. Will continue monitoring but the IOS change doesnt seem to have brought back previous
Time to Fix: